Factorial expressions
Find the value of the following:
11!11!
------
13!10!

If your answer is not an integer, write it as an exact fraction.






Answer by stanbon(57342)

Find the value of the following:
11!11!
------
13!10!
-----------
(11! / 13!) * (11! /10!)
= (1/13*12) * (11/1)
= 11/(13*12)
= 11/156
==============
Cheers,
Stan H.